OpenStudy (anonymous):

reaction: A + 2 B yields products rate data: A: 0.10 0.20 0.30 B: 0.10 0.20 0.20 Rate(M/min) 0.010 0.40 0.060 what is rate law a. k [B]^2 b. k [A] [B] c. k [A]^2 d. k [A] [B]^2

OpenStudy (aaronq):

are you sure you wrote the data correctly?

OpenStudy (anonymous):

yes but it is vertical but it is easier to write it horizontal

OpenStudy (aaronq):

Rate(M/min) 0.010 0.\(\color{red}0\)40 0.060 so it's not this ^

OpenStudy (anonymous):

o sorry no it is 0.040 I didn't c that error my bad

OpenStudy (anonymous):

0.04 = k[0.2]^n [0.2]^m 0.06 = k[0.3]^n [0.2]^m dividing these two equation gives you a simplified one of 0.667 = 0.667^n solving for n gives you "1", which is the exponent of the reactant A in the rate law "Rate = k[A]^1[B]^m", which means the reactant A follows a 1st order process

OpenStudy (anonymous):

Now we can just choose one Rate law and plug in stuff to find the order for the reactant B

OpenStudy (anonymous):

Let's choose the first experiment (column 1)

OpenStudy (anonymous):

and let's divide it with experiment 2 (column 2)

OpenStudy (anonymous):

So, 0.01 = k[0.1]^1[0.1]^m 0.04 = k[0.2]^1[0.2]^m Dividing the two equations give you a simplified one of 0.25 = 0.5 x 0.5^m 0.5 = 0.5^m so solving for m gives us 1, which is the exponent for reactant B which means B also follows 1 order kinetics

OpenStudy (anonymous):

Helloooooo anyone there?

OpenStudy (anonymous):

hi sorry yea

OpenStudy (anonymous):

So are you good with this?

OpenStudy (anonymous):

ye i am just not sure how I would do this without a calculator. It seems like this would take alot of time to do for a timed test.

OpenStudy (anonymous):

You don't need a calculator. You can make this easier to do on paper by converting these decimals to fractions and dividing or multiplying these fractions

OpenStudy (anonymous):

Ya these type of problems can be a little time consuming but if you remember the pattern of how to solve problems like these, it could take like less than a minute

OpenStudy (anonymous):

with a calculator i guess

OpenStudy (anonymous):

ok I will try to find more problems and c. thanks for the help

OpenStudy (anonymous):

np